Difference between resistance,reactance and impedence

Grade:12

1 Answers

Saurabh Koranglekar
5 years ago
Resistance is the measure of opposition to the current flow offered by the material
R = V/i

Reactance is the resistance offered to the ac currents by inductors and capacitors only. Usually denoted by X. { frequency dependent components}

forf being the frequency of AC source

For inductors X= 2πfL

For capacitors X=1/(2πfC);

These are phasor quantities

Impedance is the sum of the resistance and reactance of a circuit

denoted by Z=R+jX j is imaginary term (-1)^1/2 same as i or iota